OpenStudy (anonymous):

Solve for x. (x + 3)/(x -2) = 4

OpenStudy (anonymous):

\[\frac{ x+3 }{ x-2 }=4\]

OpenStudy (anonymous):

To get rid of the fraction, you need to multiply both sides of the equation by the denominator.

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

x = 11/3 , that's what I got...

OpenStudy (anonymous):

\[\frac{ x+3 }{ x-2 }=4\]\[x+3=4(x-2)\]\[x+3=4x-8\]\[x-4x=-8-3\]\[-3x=-11\]\[x=?\]solve it @bama42

OpenStudy (anonymous):

between that, \(\bf\huge\color{#ff0090}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) @bama42